1. Water flows into the rectangular tank through pipes A and B, with velocities of 4 m/s and 2 m/s, respectively. Determine the rate at which the surface of the water is rising. Ignore the effect of gravity on the falling water.



2. The water in the cylindrical storage tank is being drained using a pump. If the water is pumped out of the tank at 0.2 m3/s, determine the rate at which the level in the tank is descending.


1
Expert's answer
2021-12-14T09:28:40-0500

1.

"v=\frac{v_1v_2}{v_1+v_2}=2.7~\frac ms."

2.

"\dot m=\rho\dot V=200~\frac{kg}s."